# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

### v4

@@ Abstract.
The healthcare landscape is evolving, with patients seeking reliable information about their health conditions and available treatment options. Despite the abundance of information sources, the digital age overwhelms individuals with excess, often inaccurate information. Patients primarily trust medical professionals, highlighting the need for expert-endorsed health information. However, increased patient loads on experts has led to reduced communication time, impacting information sharing. To address this gap, we develop CataractBot, an experts-in-the-loop chatbot powered by LLMs, in collaboration with an eye hospital in India. CataractBot answers cataract surgery related questions instantly by querying a curated knowledge base, and provides expert-verified responses asynchronously. It has multimodal and multilingual capabilities. In an in-the-wild deployment study with 55 participants, CataractBot proved valuable, providing anytime accessibility, saving time, accommodating diverse literacy levels, alleviating power differences, and adding a privacy layer between patients and doctors. Users reported that their trust in the system was established through expert verification. Broadly, our results could inform future work on designing expert-mediated LLM bots.
@@ 1. Introduction
To address patients’ information needs, particularly during surgical treatments, we propose a doctor-integrated chatbot solution designed to provide reliable and comprehensive real-time responses. Leveraging the capabilities of large-language models (LLMs) and the widespread use of instant messaging services such as WhatsApp, this chatbot serves as an accessible, 24/7 resource capable of understanding intricate human queries and providing accurate information. In collaboration with a tertiary eye hospital in Bangalore, India, we exemplified our experts-in-the-loop chatbot approach by developing CataractBot. This WhatsApp-based chatbot, powered by LLMs, answers queries related to cataract surgery from patients and their attendants. Recognizing the limitations of generic LLMs in the medical domain, we utilized Retrieval-Augmented Generation over a knowledge base curated by doctors and hospital staff to provide hospital-specific and culturally sensitive responses. CataractBot has several features: multimodal support (accepting both speech and text inputs), multilingual capabilities (supporting five languages, including English, Hindi, and Kannada), a verification system where responses are vetted by experts (ophthalmologists for medical queries and patient coordinators for logistical ones), and a knowledge base update framework based on expert-provided edits to minimize future expert intervention. CataractBot integrates both synchronous and asynchronous messaging by providing an LLM-generated answer instantly, and subsequently providing verification at the experts’ convenience.
@@
Our work aims to answer the following research questions: (RQ1) What is the role of CataractBot in meeting the information needs of patients (and attendants) undergoing surgery? (RQ2) How do the different features of CataractBot, including experts-in-the-loop and multimodality, contribute to its usage? (RQ3) How can CataractBot be integrated into the current doctor-patient workflow? To answer these questions, we conducted an in-the-wild deployment study involving 49 information seekers111We use ‘participants’ or ‘information seekers’ to refer to patients and their attendants, and ‘experts’ to refer to doctors and coordinators. (19 patients, 30 attendants) and 6 experts (4 doctors and 2 patient coordinators). Our findings revealed that patients and attendants appreciated CataractBot because it alleviated hesitations associated with power differences in asking questions, and accommodated individuals with low literacy and tech proficiency through multilingual and multimodal support. Patients and attendants reported that their trust in the system was established through the verification performed by doctors and coordinators. Experts praised the bot as a facilitator, introducing a privacy layer between them and patients, and providing flexibility through asynchronous communication. Moreover, experts noted improved responses over time, likely due to continuous updates to the knowledge base. Our main contributions are: (1) The design and development of a novel experts-in-the-loop chatbot222Our code is publicly available at https://github.com/microsoft/BYOeB. , designed in collaboration with doctors and hospital staff, to assist patients undergoing cataract surgery with their information needs. (2) The deployment of our bot in a real-world setting, among 49 cataract surgery patients/attendants, four doctors, and two patient coordinators at a tertiary eye hospital in India. (3) Drawing from a comprehensive mixed-methods analysis of gathered data, the paper offers pivotal insights specific to CataractBot and broader implications of LLM-powered experts-in-the-loop chatbots.
@@ 2. Related Work
Our work is primarily informed by two areas of relevant research: solutions addressing patients’ information needs and the use of conversational agents, including those powered by LLMs, in healthcare settings.
@@ 2.1. Patients’ Information Needs
In summary, patients need reliable, relevant, and timely information. However, accessing this from busy doctors and hospital staff is challenging. This dichotomy results in patients either receiving generic, unverified, and hard-to-understand information from the internet and social circles or if lucky, obtaining personalized, verified answers, albeit at a higher cost, from healthcare experts. Our solution aims to strike a balance by providing a generic, high-quality answer instantly using LLMs, which is later verified by a medical professional with minimal increase to their workload.
2.2. Chatbots in Healthcare
Chatbots enable natural language conversation, resulting in a minimal learning curve and offering a personalized experience (Jain et al., 2018; Jo et al., 2023; Roller et al., 2020; Ferré et al., 2020; Chung, 2017; Desai and Chin, 2023). They enable users to build complex queries message by message while retaining conversation context (Xiao et al., 2023). Researchers have applied chatbots across diverse healthcare settings (Wutz et al., 2023), including scheduling appointments (Ittarat et al., 2023), collecting pre-consultation information (Li et al., 2024), assisting new mothers (Yadav et al., 2019), supporting individuals suffering from chronic diseases (Montagna et al., 2023; Hao et al., 2024), and offering mental health support (Kumar et al., 2023; Jo et al., 2023; Ma et al., 2023; Kim et al., 2024). Commercial chatbots, such as Babylon Health, Ada, and Florence333Babylon Health: https://www.emed.com/uk. Ada: https://ada.com/. Florence: https://florence.chat/., allow individuals to inquire about their health issues. All of these bots are fully automated, with some even using LLMs. However, the involvement of doctors in these bots is limited to the data curation step (Xiao et al., 2023). Consequently, they are susceptible to generating inaccurate information, rendering them “not ready for clinical use” (Au Yeung et al., 2023).
LLMs have been criticized for hallucinating and not displaying the reason behind their outputs (Fogel and Kvedar, 2018). Several methods have been proposed to address these issues, including chain-of-thought prompting (Wei et al., 2024), retrieval-augmented generation (Lewis et al., 2020), and few-shot learning. Prior work suggests that healthcare professionals are more likely to adopt AI tools when they perceive them as ‘partners’ that enhance their workflow and expertise (Henry et al., 2022). Seitz et al. (2022) identified that patients’ trust in healthcare chatbots is primarily based on a rational evaluation of their capabilities, while trust in human medical professionals relies more on emotional connections and personal rapport. Integrating doctors into the loop with automated chatbots could potentially combine the best of both worlds, a concept we propose in CataractBot.
Given our specific focus on cataract surgery, we also examined literature at the intersection of ophthalmology and LLMs. Bernstein et al. (2023) reported that responses generated by LLMs to patient eye-care-related questions are comparable to those written by ophthalmologists. Based on similar findings, Ittarat et al. (2023) proposed integrating chatbots into ophthalmology practices to provide 24/7 support for common inquiries on eye conditions. Additionally, Yilmaz and Dogan (2024) reported that chatbots are an effective tool for educating cataract surgery patients. Our work contributes to this body of literature.
@@ 2.2. Chatbots in Healthcare
Beyond chatbots, prior work has explored “chat” as an interface to connect patients and doctors in both synchronous (Li et al., 2023) and asynchronous (Johansson et al., 2020a, b) manners. Synchronous messaging requires both the patient and the doctor to be simultaneously active, allowing instant feedback and coherent dialogue. On the other hand, asynchronous messaging platforms like WhatsApp (Yadav et al., 2019), WeChat (Wang et al., 2020), SMS (Perrier et al., 2015), email (Makarem and Antoun, 2016), or web portals (Mamykina et al., 2008; Tuli et al., 2018) enable communication at the convenience of both parties, providing flexibility. Previous research indicates that both doctors and patients prefer asynchronous texts over instant audio/video calls (Stamenova et al., 2020; Johansson et al., 2020a), despite lacking instant feedback (Li et al., 2023). In our work, we combine both approaches—leveraging LLM-based responses in the synchronous mode and having a human medical expert respond in the asynchronous mode. This hybrid approach represents a novel and effective method for addressing information needs in the healthcare domain.
@@
Finally, real-world healthcare deployments of AI are limited due to socio-technical uncertainties in the last mile (Osman Andersen et al., 2021). To avoid such shortcomings, as recommended by Thieme et al. (2023b, a), we designed and studied our system within a tertiary eye hospital in India and understood its integration into clinical workflows.
@@ 3. Formative Study
The development of CataractBot was informed by our literature survey, our understanding of the Indian healthcare ecosystem, and insights from a formative study conducted at Sankara Eye Hospital. To identify requirements, we conducted semi-structured interviews with 4 ophthalmologists and 2 patient coordinators between July-Aug 2023. From these interviews, we aimed to understand current workflows, the types of questions patients ask, and their concerns regarding patients’ knowledge gaps. Participants had an average age of 38.58.8 years, with 11.07.5 years of work experience. They participated in the study voluntarily without compensation. The interviews were conducted in English and lasted around 45 minutes each. All sessions were audio-recorded and transcribed by the interviewer. We conducted thematic analysis to identify key design requirements for the chatbot.
@@
Accuracy: The chatbot’s ability to answer both medical and logistical questions with accuracy, precision and contextual relevance was found to be essential given the medical context. While a standard LLM like GPT-4 (OpenAI et al., 2023) might offer accurate information, it may not be tailored to India- or hospital-specific nuances. For instance, doctors informed us that although the phacoemulsification method is prevalent globally for cataract surgeries, they often favor manual small incision-based procedures (Singh et al., 2017) due to their cost-effectiveness. Similarly, cultural considerations significantly influence pre- and post-surgery questions. As a doctor stated, “People will have weird questions… like ‘when can I do ayurvedic nasal therapy post-op?’.” Questions related to diet require awareness of Indian food options for accurate guidance. To ensure accurate and contextual responses, our chatbot relies on a knowledge base curated by hospital medical staff, including doctors, patient coordinators, and members of Quality Control and Patient Safety team.
@@
Trustworthiness: Access to trustworthy information is crucial during medical treatment. As one doctor noted, “The success of treatment is not only in medicine, but the trust they place on us.” Patients and their attendants typically place their utmost trust in the operating doctor and hospital staff. Even when patients learn something from the internet or other sources, they often confirm that information with their doctor. E.g., a doctor noted, “They come and ask us… my friends have told me this… is it correct or not?” To ensure trustworthiness in CataractBot we decided that (a) each response will be verified and, if necessary, edited by doctors and patient coordinators, and (b) patients (or their attendants) will be explicitly notified when this expert verification happens. Note: We intentionally designed CataractBot to be non-specific to individual patients by not connecting it with the hospital’s scheduling system or EHR data. This decision was made because LLMs can memorize training data, including personally identifiable information (PII) like emails and phone numbers, and leak it during inference (Borkar, 2023), raising privacy concerns and potentially breaking users’ trust.
@@
Timeliness: Discussions with hospital staff highlighted the crucial need for real-time responses. Patient coordinators noted that patients frequently reach out to them over multiple phone calls to ask logistical questions and relay medical queries to their consulting doctors. Due to busy schedules of doctors and coordinators, immediate responses are not always feasible. This often results in patient seeking information from other sources (including friends and online sources), which can propagate misinformation, or fail to completely address their information needs. To counter this, we decided that the chatbot should provide an instantaneous response to every patient query using a custom knowledge base curated by hospital staff. Subsequently, these responses get reviewed and refined by doctors/coordinators. Moreover, doctors mentioned their unavailability for several hours while performing surgeries. o ensure swift verification of every query, CataractBot forwards queries to an ‘escalation’ expert if they remains unverified for 3 hours.
@@
Usability: After cataract surgery, patients can reference their discharge summaries for any queries, but this is challenging due to their temporarily reduced vision. Prior research (Grönvall and Verdezoto, 2013) has identified the crucial role family members play in monitoring the patient’s health. Coupled with doctors’ input (“Patients ask hardly 10% of the question… attendants are the ones who ask.”), this underscored the necessity for the bot to be usable by four key stakeholders of the cataract surgery ecosystem: patients, their attendants, doctors, and patient coordinators. To cater to diverse backgrounds, literacy levels, and varying degrees of tech-literacy among hospital visitors, as suggested by doctors and patient coordinators, we made the chatbot multilingual and multimodal, similar to prior work (Jain et al., 2018; Findlater et al., 2009). We chose WhatsApp as the messaging platform due to its widespread use in India (Statista, 2022) and therefore minimal learning curve and quick onboarding without additional installations, despite the richer feature set of other platforms like Telegram or the development flexibility of a custom chatbot. Considering doctors’ demanding schedules, often involving consulting 50 patients and performing 10 cataract surgeries a day, we implemented features to minimize their bot-related workload: (1) one-click interaction to verify an answer, (2) allow experts to provide corrections using informal and ‘symbolic’ messages (Beasley, 2009), (3) handle spelling and grammar errors, and (4) use expert edits to enhance the bot’s knowledge base, minimizing similar edits in future.

### v5

@@ Abstract.
The healthcare landscape is evolving, with patients seeking reliable information about their health conditions and available treatment options. Despite the abundance of information sources, the digital age overwhelms individuals with excess, often inaccurate information. Patients primarily trust medical professionals, highlighting the need for expert-endorsed health information. However, increased patient loads on experts has led to reduced communication time, impacting information sharing. To address this gap, we developed CataractBot111Our code is publicly available at https://github.com/microsoft/BYOeB.. CataractBot answers cataract surgery related questions instantly using an LLM to query a curated knowledge base, and provides expert-verified responses asynchronously. It has multimodal and multilingual capabilities. In an in-the-wild deployment study with 49 patients and attendants, 4 doctors, and 2 patient coordinators, CataractBot demonstrated potential, providing anytime accessibility, saving time, accommodating diverse literacy levels, alleviating power differences, and adding a privacy layer between patients and doctors. Users reported that their trust in the system was established through expert verification. Broadly, our results could inform future work on designing expert-mediated LLM bots.
@@ 1. Introduction
To address patients’ information needs, particularly during surgical treatments, we propose a chatbot solution integrated with medical professionals to provide accessible, reliable and comprehensive responses. Leveraging the capabilities of large-language models (LLMs) and the widespread use of smartphones and instant messaging services such as WhatsApp, this chatbot serves as a 24/7 resource capable of understanding intricate human queries and providing accurate information. In collaboration with a tertiary eye hospital in Bangalore, India, we exemplified this experts-in-the-loop chatbot approach by developing CataractBot, which answers queries related to cataract surgery from patients and their attendants. Recognizing the limitations of generic LLMs in the medical domain, we utilized Retrieval-Augmented Generation (Lewis et al., 2020) with an LLM over a custom knowledge base curated by doctors and hospital staff to provide hospital-specific and culturally sensitive responses. CataractBot has several features to enhance adoption and sustained use. Capitalizing on the widespread adoption of smartphones, we chose a mobile-first design paradigm, deploying the chatbot entirely on the ubiquitous WhatsApp platform. To cater to diverse hospital visitors with varying literacy levels and technological proficiency, we made CataractBot multilingual (supporting five languages, including English, Hindi, and Kannada) and multimodal (accepting both speech and text inputs). We use an LLM to process complex and ill-formed queries, generating instant responses from the curated knowledge base. Crucially, CataractBot incorporates a system where the LLM’s responses are asynchronously verified by experts and corrected if needed. Ophthalmologists verify answers to medical queries and patient coordinators to logistical ones. These expert-provided edits are used to update the knowledge base, to minimize future expert intervention.
@@
Our work aims to answer the following research questions: (RQ1) What is the role of CataractBot in meeting the information needs of patients (and attendants) undergoing surgery? (RQ2) How do the different features of CataractBot, including LLM-generated answers, experts-in-the-loop and multimodality, contribute to its usage? (RQ3) How can CataractBot be integrated into the current doctor-patient workflow? To answer these questions, we conducted an in-the-wild deployment study involving multiple stakeholders (See Table 1): 49 information seekers (19 patients, 30 attendants) and 6 experts (4 doctors and 2 patient coordinators). Our findings indicated that patients and attendants appreciated CataractBot because it alleviated hesitations associated with power differences in asking questions, and accommodated individuals with low literacy and tech proficiency through multilingual and multimodal support. Patients and attendants reported that their trust in the system was established through the verification performed by doctors and coordinators. Experts praised the bot as a facilitator, introducing a privacy layer between them and patients, and providing flexibility through asynchronous communication.
Our main contributions are: (1) The design and development of a novel LLM-powered, experts-in-the-loop, multilingual, multimodal chatbot, created in collaboration with doctors and hospital staff to assist cataract surgery patients with their information needs. (2) The deployment of our bot in a real-world multi-stakeholder setting, among 49 cataract surgery patients/attendants, four doctors, and two patient coordinators at a tertiary eye hospital in India. (3) Drawing from a comprehensive mixed-methods analysis of gathered data, the paper offers pivotal insights specific to CataractBot and broader implications of LLM-powered experts-in-the-loop chatbots.
@@ 2. Related Work
Our work is primarily informed by two areas of relevant research: solutions addressing patients’ information needs and the use of conversational agents, particularly those powered by LLMs, in healthcare settings.
@@ 2.1. Patients’ Information Needs
Prior work has also explored “chat” as an interface to address patients’ information needs, by connecting them with doctors in both synchronous (Li et al., 2023) and asynchronous (Johansson et al., 2020a, b) manners. Synchronous messaging requires both patient and doctor to be simultaneously active, allowing instant feedback and coherent dialogue. On the other hand, asynchronous messaging platforms like WhatsApp (Yadav et al., 2019), WeChat (Wang et al., 2020), SMS (Perrier et al., 2015), email (Makarem and Antoun, 2016), or web portals (Mamykina et al., 2008; Tuli et al., 2018) enable communication at the convenience of both parties. Research indicates that both doctors and patients prefer asynchronous texts over instant audio/video calls (Stamenova et al., 2020; Johansson et al., 2020a), despite lacking instant feedback (Li et al., 2023).
In summary, patients need reliable, relevant, and timely information. However, obtaining this from busy doctors and hospital staff is challenging. This dichotomy results in patients either receiving generic, unverified, and hard-to-understand information from the internet and social circles, or, if lucky, obtaining personalized, verified answers from healthcare experts—at a higher cost. Our solution aims to strike a balance by providing a generic, high-quality LLM answer, later verified by a medical professional with minimal increase to their workload.
2.2. LLM Chatbots in Healthcare
Chatbots enable natural language conversations, offering a minimal learning curve and a personalized experience (Thirunavukarasu et al., 2023). AI-powered chatbots surpass rule-based platforms in language understanding, allowing users to build complex queries message by message while retaining context (Xiao et al., 2023). With the advent of LLMs, building complex chatbots has become significantly easier, enabling researchers to apply LLM-powered bots across diverse healthcare settings (Thirunavukarasu et al., 2023), including pre-consultation data collection (Li et al., 2024), chronic disease management (Montagna et al., 2023; Hao et al., 2024), and mental health support (Kumar et al., 2023; Jo et al., 2023; Kim et al., 2024). Commercial solutions, such as Babylon Health (eMed, 2024), now provide individuals with instant access to any health-related information. However, LLM-powered healthcare tools have generally been criticized for hallucinations, errors, and a lack of transparency in their reasoning (Au Yeung et al., 2023; Denecke et al., 2024). As the involvement of doctors in these bots is–at most–limited to the data curation step (Xiao et al., 2023), they are susceptible to generating inaccurate information and are “not ready for clinical use” (Au Yeung et al., 2023).
Several methods have been proposed to address these issues, including chain-of-thought prompting (Wei et al., 2024), retrieval-augmented generation (Lewis et al., 2020), and few-shot learning. Seitz et al. (2022) identified that patients’ trust in healthcare chatbots is primarily based on a rational evaluation of their capabilities, while trust in human medical professionals relies more on emotional connections and personal rapport. Integrating doctors into the loop with automated LLM chatbots could combine the best of both worlds, a concept we propose in CataractBot. In our work, we leverage LLM-based responses in real-time, while human medical experts verify and respond asynchronously. This hybrid approach is a novel and reliable method for addressing healthcare information needs.
We chose cataract surgery–one of the world’s most common procedures (McGhee et al., 2020)–as the use case for developing and testing our system. This decision was motivated by our longstanding collaboration with Sankara Eye Hospital, an eye hospital in India. Also, cataract surgery involves well-defined protocols and common patient queries, making it an ideal starting point for testing the bot’s ability to deliver accurate, standardized information from a curated knowledge base. However, our solution is versatile, and we believe it can effectively support a range of health conditions, including mental health and chronic care. Given our specific focus on cataract surgery, we examined literature at the intersection of ophthalmology and LLMs. Bernstein et al. (2023) reported that responses generated by LLMs to patient eye-care-related questions are comparable to those written by ophthalmologists. Based on similar findings, Ittarat et al. (2023) proposed integrating chatbots into ophthalmology practices to provide 24/7 support for common inquiries on eye conditions. Additionally, Yilmaz and Dogan (2024) reported that chatbots are an effective tool for educating cataract surgery patients. Our work contributes to this body of literature.
@@ 2.2. LLM Chatbots in Healthcare
Finally, real-world healthcare deployments of AI are limited due to socio-technical uncertainties in the last mile (Osman Andersen et al., 2021). To avoid such shortcomings, as recommended by Thieme et al. (2023b, a), we designed and studied our system within a tertiary eye hospital in India and examined its integration into clinical workflows.
@@ 3. Formative Study
The development of CataractBot was informed by our literature survey, our understanding of the Indian healthcare ecosystem, and insights from a formative study conducted at Sankara Eye Hospital. To identify requirements, we conducted semi-structured interviews with 4 ophthalmologists and 2 patient coordinators between July-Aug 2023. We solely interviewed experts, as we aimed to understand the status quo, including their workflows, the types of questions patients ask, and experts’ concerns regarding patients’ knowledge gaps. We then conducted pilot tests with all end users–patients, attendants, doctors, and coordinators–which iteratively informed the design of CataractBot. Interview participants had an average age of 38.28.3 years, with 11.37.3 years of work experience (Table 4 in Appendix A.2). They participated in the study voluntarily without compensation. The interviews were conducted in English and lasted around 45 minutes each. All sessions were audio-recorded and transcribed by the interviewer. The interviewer open coded the interview transcripts, following inductive thematic analysis (Braun and Clarke, 2006) to identify key design requirements for the chatbot. Two researchers met regularly to review the emerging codes and finalized the high-level themes.
3.1. Background
Our design requirements were shaped by the workflows and information needs specific to cataract surgery at Sankara Eye Hospital. In the current workflow, patients (with their attendants) first consult a doctor who diagnoses the cataract’s maturity and type, recommending the urgency of surgery and potential complications. After deciding on surgery, the patient and attendant meet with a patient coordinator for guidance on pre- and post-operative measures, lens options across different budget ranges, and logistical tasks such as insurance and scheduling. Patients also undergo pre-operative medical tests. For further questions, they can either visit the hospital to meet the doctor or contact the coordinator by phone. On the day of surgery, patients arrive with their attendants in the morning, undergo surgery by mid-morning, and are discharged by the afternoon or evening. The doctor and coordinator provide post-surgery care instructions, covering medication, exercise, bathing, travel, and screen time. One week post-surgery, patients return for a follow-up appointment with the operating doctor.
3.2. Design Requirements
@@
Accuracy. The chatbot’s ability to answer both medical and logistical questions with accuracy, precision and contextual relevance was found to be essential given the medical context. While a standard LLM like GPT-4 (OpenAI et al., 2023) might offer accurate information, it may not be tailored to India- or hospital-specific nuances. For instance, doctors informed us that although the phacoemulsification method is prevalent globally for cataract surgeries, they often favor manual small incision-based procedures (Singh et al., 2017) due to their cost-effectiveness. Similarly, cultural considerations significantly influence pre- and post-surgery questions. As a doctor stated, “People have unique questions… like ‘when can I do nasal yoga post-op?’.” Further, queries related to diet require awareness of Indian food options for precise guidance. Thus, our bot relies on a custom knowledge base curated by hospital staff.
@@
Trustworthiness. Access to trustworthy information is crucial during medical treatment. As one doctor noted, “The success of treatment is not only in medicine, but the trust they place on us.” Patients and their attendants typically place their utmost trust in the operating doctor and hospital staff. Even when patients learn something from the internet or other sources, they often confirm that information with their doctor. To ensure trustworthiness, we decided that (a) each response would be verified and, if necessary, edited by a doctor or patient coordinator, and (b) patients (or their attendants) would be explicitly notified upon expert verification. Note: We intentionally designed CataractBot to be non-specific to individual patients by not connecting it with the hospital’s scheduling system or EHR data. This decision was made because LLMs can memorize training data, including personally identifiable information (PII) like emails and phone numbers, and leak it during inference (Borkar, 2023), raising privacy concerns and potentially breaking users’ trust.
@@
Timeliness. Discussions with hospital staff highlighted the crucial need for real-time responses. Patient coordinators noted that patients frequently reach out to them over multiple phone calls to ask logistical questions and relay medical queries to their consulting doctors. Due to experts’ busy schedules, immediate responses are not always feasible. This often results in patient seeking information from other sources (including the internet), which can propagate misinformation, or fail to completely address their information needs. To counter this, we decided that the chatbot should provide an instantaneous response to every patient query using the LLM and a custom knowledge base. Subsequently, these responses get reviewed and refined by doctors/coordinators. Moreover, doctors mentioned their unavailability for several hours while performing surgeries. To ensure swift verification, CataractBot forwards queries to an ‘escalation’ expert if they remains unverified for 3 hours.
@@
Usability. After cataract surgery, patients can reference their discharge summaries for any queries, but this is challenging due to their temporarily reduced vision. Prior research (Grönvall and Verdezoto, 2013) has identified the crucial role family members play in monitoring the patient’s health. Coupled with doctors’ input (“Patients ask hardly 10% of the question… attendants are the ones who ask.”), this underscored the necessity for the bot to be usable by both the patients and their attendants. To cater to diverse backgrounds, we made the chatbot multilingual and multimodal. We chose WhatsApp as the messaging platform due to its widespread use in India (Statista, 2022) and therefore minimal learning curve and quick onboarding without additional installations, despite the richer feature set of other platforms like Telegram or the development flexibility of a custom chatbot. Considering doctors’ demanding schedules, often involving consulting 50 patients and performing 10 cataract surgeries a day, we implemented features to minimize their bot-related workload: (1) one-click interaction to verify an answer, (2) allow experts to provide corrections using informal messages (Beasley, 2009), (3) handle spelling and grammar errors, and (4) use expert edits to enhance the bot’s knowledge base, minimizing similar edits in future.